OpenStudy (anonymous):

Is there a method other than trial and error to solve Asin(x) +Bcos(x) = C ?

OpenStudy (kira_yamato):

Use the R-formula.... a cos x + b sin x = R cos (x - Ω) Where \[R = \sqrt{a^2 + b^2}\] and \[\tan Ω = b/a\]